During 2018 and 2019, Kale Co. completed the following transactions relating to its bond issue. The companys fiscal year endDuring 2018 and 2019, Kale Co. completed the following transactions relating to its bond issue. The companys fiscal year endDuring 2018 and 2019, Kale Co. completed the following transactions relating to its bond issue. The companys fiscal year end
Answer #1

A.
Market rate of interest was more than the stated rate of interest, since bonds are issued at discount
Cash Received = $348000

B.
Discount on issue of bonds = $360000 - $348000 = $12000
Amortization of Discount for every interest = $12000/16 = $750

2018 2019
Liabilities
Interest Payable $              9,600 $                9,600
Bonds Payable $         3,60,000 $          3,60,000
Less : Discount on Bonds Payable $            10,750 $                9,250
Carrying Value of Bonds Payable $         3,49,250 $          3,50,750
Total Liabilities $        3,58,850 $          3,60,350

Calculation

2018 2019
Liabilities
Interest Payable =360000*8%*4/12 9600
Bonds Payable 360000 360000
Less : Discount on Bonds Payable =12000-1500*10/12 =10750-1500
Carrying Value of Bonds Payable
Total Liabilities

C & D

2018 2019
Interest Expense $            25,250 $              30,300
Interest Paid $            14,400 $              28,800
2018 2019
Interest Expense =360000*8%*10/12+1500*10/12 =360000*8%+1500
Interest Paid =360000*8%*6/12 =360000*8%